OpenStudy (anonymous):

how do i solve for x the equation is 4x^+28x=176

OpenStudy (anonymous):

Take 176 to the left side, 4x^2 + 28x - 176 = 0 Divide all the terms by 4 x^2 + 7x - 44 = 0 Then factor.

OpenStudy (anonymous):

^what

OpenStudy (anonymous):

subtract 176 from both sides, and then use the quadratic formula

OpenStudy (anonymous):

whats the power

OpenStudy (anonymous):

that is a lengthy one 11,4 as 11*-4=-44& 11-4=7 so x=11,-4:)

OpenStudy (anonymous):

4x(squared) + 28x = 176

OpenStudy (anonymous):

aperogalics--can you show me the work so I can understand the solution?

OpenStudy (anonymous):

simple if the factorized form is(x-a)(x-b)=x^2-[a+b]x+ab i just use hit and trial method to find 2 nos whose sum or difference(depends upon product sign) is 7 and product is -44:)

OpenStudy (anonymous):

4x^2 + 28x = 176 4x^2 + 28x - 176 = 0 4 (x^2 + 7x - 44) = 0 4 (x + 11) (x - 4) = 0 ignore the 4 out front and you get two equations, x + 11 = 0 and x - 4 = 0 therefore, x = -11 and x = 4

hero (hero):

isellsani, all you have to do is divide both sides by 4, then you would have (x + 11)(x - 4) = 0

OpenStudy (anonymous):

I know that, was just avoiding another step as it's mundane.

hero (hero):

It's not a mundane step, it is simply the next logical one.